OpenStudy (anonymous):

Which ratio is equivalent to 6 : 7? A.12 : 13 B.18 : 48 C.36 : 49 D.48 : 56

OpenStudy (jojo4eva):

D. 48:56

OpenStudy (jojo4eva):

got it??

OpenStudy (anonymous):

yea @jojo4eva

OpenStudy (anonymous):

it is just like when you do 2x=4 what you do to 1 side you do to the other. 6:7 multiply both sides by 8 and you get 48:56

OpenStudy (anonymous):

D because 6*8=48 and 7*8=56 so since the numerator and denominator were both multiplied by the same number, 8, the ratio 48/56 is equivalent to 6/7

OpenStudy (anonymous):

Which ratio is equivalent to 12 : 32?

OpenStudy (anonymous):

an easy way to do this for your test! pick one of them like 12:13 divide the 12 by the 6 because they are both on the left side. you get 2. so if 7*2 = 13 that would be right. (it obviously doesnt)